OpenStudy (anonymous):

Where are the asymptotes of f(x) = tan 2x from x = 0 to x = π?

OpenStudy (anonymous):

you have asymptotes because the function of tan is undefined at pi/2 and 3pi/2

OpenStudy (anonymous):

your domain is restricted to pi so it is pi/2

OpenStudy (anonymous):

ok thanks!

OpenStudy (mathmate):

@hotguy hint: Tan(x) has asymptotes whenever x=n\(\pi\)/2, where n=1,3,5,... So when f(x)=tan(2x) for x=0 to pi, there are two values of x (0\(\le\)x\(\le \pi\) ) to make f(x) = tan(pi/2) and f(x)=tan(3pi/2) see if you can find the values of x.